OpenStudy (bowtie_man123):

Which is a solution to the equation y = 2x – 5? A. (5, –5) B. (3, 11) C. (2, 1) D. (1, –3)

OpenStudy (bowtie_man123):

Thanks and i cant remember how to do this would you be able to give me an explanation?

OpenStudy (eli_moses):

ok so basically i just simplified the coordinates into the equation for x and y

OpenStudy (bowtie_man123):

OHHHHHHHHHHHH ok thanks

OpenStudy (eli_moses):



OpenStudy (eli_moses):

you can simplify it from there

OpenStudy (tkhunny):

It is a substitution problem. You have this: y = 2x – 5 For part A, you have this: (5, –5) ==> Substitute x = 5 and y = -5 Thus: y = 2x – 5 ==> -5 = 2(5) - 5 OR -5 = 10-5 OR -5 = 5 That last result clearly is incorrect, so you are done. (5,-5) is NOT a solution.

OpenStudy (bowtie_man123):

what about this one Which is not a solution to the equation ? A. (–4, –7) B. (3, 14) C. (2, –1) D. (1, 8)

OpenStudy (eli_moses):

is it the same equation?

OpenStudy (tkhunny):

It is EXACTLY the same. Show YOUR work. Do NOT post any more problems with no work shown.